OpenStudy (tyrinzoebrown):

|x|+5=18 5 or -5 13 or -13 18 or -18 23 or -23

OpenStudy (anonymous):

|x|+5-5=18-5 |x|=13 x can be 13 or -13

OpenStudy (tyrinzoebrown):

I thought that was the answer but that negative didnt make sence to me thank you

OpenStudy (anonymous):

absolute value refers to distance so anything inside the absolute value signs is positive no matter what

OpenStudy (anonymous):

are the numbers above choices for the answers

OpenStudy (anonymous):

5 or -5 choices? 13 or -13 18 or -18 23 or -23